OpenStudy (anonymous):

Solve Each Equation by first clearing the fractions or decimals 0.04n-0.05(n+2)=0.1 I Need to show all work. The answer is -20

OpenStudy (anonymous):

\[0.04n-0.05(n+2)=0.1 *100 4n-5(n+2)=10 4n-5n-10=10 -n=20 n=-20\]

OpenStudy (anonymous):

\[100 (0.04n - 0.05 (n+2) = 100 (0.04n - 0.05n-0.10)\]\[4n - 5n - 10 = -n -10\]\[-n-10 = -20 \]\[-n = -10\]\[n=10\]